Cobb Doglas production function is given by y=10(LK)1/2.find the expression for the total differential.


1
Expert's answer
2021-05-30T14:15:28-0400

 Differentials of output with respect to labor and capital are the marginal product of capital and marginal product of labor respectively. The differential represents a small change.

"y=10(LK)^{0.5}"

"\frac{\delta y}{\delta K}=10\sqrt{L}(\frac{1}{2\sqrt{K}})"

"=5\sqrt{\frac{L}{K}}"


"y=10\sqrt{LK}"

"\frac{\delta y}{\delta L}=10\sqrt{K}(\frac{1}{2\sqrt{L}})"

"=5\sqrt{\frac{K}{L}}"